# On the third homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$

### v1

@@ 4. A refined Bloch-Wigner exact sequence
We actually can calculate the map
$\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
which appears in the above theorem. Let $A$ be a local domain such that ${\rm char}(A)\neq 2$. Let $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$. Then by some tedious calculation one can show that under the surjective map $H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)$, the cycle class of
$[a|b]+R_{ab}-R_{a}-R_{b}+R_{1}\in B_{2}(\mathit{{\rm SL}}_{2}(A))_{\mathit{{\rm SL}}_{2}(A)}$
maps to $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$, where for any $z\in{A^{\times}}$,
$R_{z}$ $=[w|g^{-1}_{z}]-[w|h^{-1}_{z}]-[h^{-1}_{z}D(z)^{-1}|wh^{-1}_{z}]+[h_{z^{-1}}|w^{-1}]-[h^{-1}_{z}|D(z)^{-1}]$
$+[D(z)h_{z}|D(z)^{-1}]-[D(z)h_{z}|h^{-1}_{z}]+2[h^{-1}_{z}|h^{-1}_{z}]+[h^{-2}_{z}|h^{-2}_{z}]$
$+[D(2)h^{-1}_{z}D(z)^{-1}|D(2)]-[D(2)|h^{-1}_{z}],$
with $g_{z}=\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!z\end{array}\!\!\right)$ and $h_{z}=\left(\!\!\begin{array}[]{cc}\!\!1&\!\!z^{-1}\\ \!\!0&\!\!1\end{array}\!\!\right)$. For another cycle that its class maps to $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$ see [15, §7]. Now $(-1)\otimes\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle$ maps in $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ to the class of the cycle which is obtained from the shuffle product of $[-1]\in B_{1}(\mu_{2}(A))_{\mu_{2}(A)}$ with the cycle $[a|b]+R_{ab}-R_{a}-R_{b}+R_{1}$.
@@
If $F$ is a real closed field, then the sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{RB}(F)\rightarrow 0$
is exact.
Proof.
It is well-know that for any real closed field $F$, $I(F)\simeq\mathbb{Z}$ which is generated by $\langle\!\langle-1\rangle\!\rangle$ and $I^{2}(F)=2I(F)\simeq 2\mathbb{Z}$. Consider the exact sequence
[…]
(discussed before Theorem 4.3). Note that
$I(F)\otimes_{\mathbb{Z}}\mu_{2}(F)\simeq I(F)/2I(F)\simeq I(F)/I^{2}(F)\simeq\mathcal{G}_{F}\simeq\mathbb{Z}/2.$
Applying the Snake lemma to the commutative diagram
[…]
we obtain the exact sequence
[…]
We show that ${E^{\prime}}_{1,2}^{2}=0$. Consider the differential
${d^{\prime}}_{2,2}^{1}:H_{2}(\mathit{{\rm SL}}_{2}(F),Z_{1}(F^{2}))\rightarrow H_{2}(T(A),\mathbb{Z})\simeq F^{\times}\wedge F^{\times}=F^{>0}\wedge F^{>0}$
(note that $F=\mu_{2}(F)\oplus F^{>0}$ and $F^{>0}=(F^{>0})^{2}$). For any $a,b\in F^{>0}$, let
$Y_{a,b}:=\overline{([D(a)|D(b)]-[D(b)|D(a]))\otimes\bigg{(}(\langle{\boldsymbol{e}_{1}}\rangle,\langle{\boldsymbol{e}_{2}}\rangle)+(\langle{\boldsymbol{e}_{2}}\rangle,\langle{\boldsymbol{e}_{1}}\rangle)\bigg{)}}\in H_{2}(\mathit{{\rm SL}}_{2}(F),Z_{1}(F^{2})).$
It is straightforward to check that
${d^{\prime}}_{2,2}^{1}(Y_{a,b})=2(a\wedge b).$
Since $F^{>0}=(F^{>0})^{2}$, ${d^{\prime}}_{2,2}^{1}$ is surjective and thus ${E^{\prime}}_{1,2}^{2}=0$. Moreover, it is not difficult to show that ${E^{\prime}}_{0,3}^{2}\simeq(\bigwedge_{\mathbb{Z}}^{3}F^{\times})/2\oplus{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))={{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))$ (see [23, Proposition 2.7] and the paragraph after that). Now by an easy analysis of the spectral sequence ${E^{\prime}}_{\bullet,\bullet}^{1}$, we obtain the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow{E^{\prime}}^{\infty}_{2,1}\rightarrow 0$
From the commutative diagram with exact rows
[…]
it follows that the middle vertical map has a kernel with at most four elements. But we have seen that the composite map
$\mathbb{Z}/4\simeq H_{3}(\langle w\rangle,\mathbb{Z})\hookrightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{RB}(F)$
is trivial (see the proof of Theorem 2.1). This shows that the kernel of the surjective map $H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\twoheadrightarrow\mathcal{RB}(F)$ has exactly four elements. Hence we are done. ∎
Corollary 4.7.
If $F$ is a real closed field, then the sequence
$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\rightarrow 0$
@@
This follows from Proposition 4.6, Theorem 4.3 and Corollary 2.4. ∎

### v3

@@ 3. Low dimensional homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
The result of the above proposition is valid for any domain $A$ such that $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is torsion of odd exponent. For example let $\mathcal{O}_{-3}$ be the ring of algebraic integers of the imaginary quadratic field $\mathbb{Q}(\sqrt{-3})$. The homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ have been calculated in [2, Theorem, page 376] and [25, Theorem 5.7]. We list these homology groups for $1\leq n\leq 3$:
From Theorem 3.1 we have the exact sequence
is non-trivial.
Example 5.3.
Let $A$ be a local domain such that $A/\mathfrak{m}_{A}\simeq\mathbb{F}_{2}$. Then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq A/\mathfrak{m}_{A}^{2}$. By a careful analysis of the composition
Observe that
If $A=\mathbb{Z}_{(2)}=\{a/b\in\mathbb{Q}:a,b\in\mathbb{Z},2\nmid b\}$ (or $A=\mathbb{Z}_{2}$, the ring of $2$-adic integers), then $2\notin\mathfrak{m}_{A}^{2}$ (by the Nakayama lemma). Thus the sequence
is exact. Note that $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq\mathbb{Z}/4$ [19, Example 4.2]. With an argument similar to the proof of Proposition 4.1(ii), one can show that
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(\mathbb{Z}_{(2)}),H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z}))\rightarrow\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}$
is trivial and
$H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq\mathbb{Z}/2$
is surjective. Thus
$H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq H_{2}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})$
and the sequence
is exact. We have similar results for $\mathbb{Z}_{2}$. Hence Proposition 5.1 is not necessarily true when $|A/\mathfrak{m}_{A}|=2$.
@@
Let $F$ be either a finite field, a quadratically closed field, or a real closed field. Then the sequence
@@
(i) Let $F=\mathbb{F}_{q}$ be finite. For fields of characteristic $2$ the above exact sequence is trivial. Thus we may assume that $q$ is odd. We have seen that
$H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\simeq\begin{cases}0&\text{if $q\neq 3$}\\ \mathbb{Z}/3&\text{if $q=3$.}\end{cases}$
Moreover, it is well-known that
$H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\simeq\begin{cases}0&\text{if $q\neq 9$}\\ \mathbb{Z}/3&\text{if $q=9$}\end{cases}$
(see [24, page 280]). These facts together with Theorem 3.1 give us the desired exact sequence.
(ii) Let $F$ be quadratically closed. By Proposition 2.2, $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})$ is uniquely 2-divisible. Thus by Proposition 5.1 we obtain the desired exact sequence.
(iii) Let $F$ be a real closed field. By Proposition 2.4 we have the Bloch-Wigner exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{B}(F)\rightarrow 0$
From [20, Corollary 3.5] and [7, Corollary 4.7] we obtain the isomorphism
$H_{3}(\mathit{{\rm PSL}}_{2}(F),\mathbb{Z})\simeq\mathcal{B}(F).$
Our claim follows from these two facts. Thus we presented a proof of this in above proposition. ∎
Remark 5.5.
It is mentioned in [11, page 230] that the sequence
$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{R}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{R}),\mathbb{Z})\rightarrow 0$
is exact and for a proof the authors referred to [10, App. A]. We couldn’t really verify this. So we gave a proof of this in above proposition.